The present age of a father is four times that of his son, in five years the father's age will be three times that of his son.find the present ages of son and father by simultaneous equation
Answer by stanbon(75887)

The present age of a father is four times that of his son, in five years the father's age will be three times that of his son.find the present ages of son and father by simultaneous equation
-------
f = 4s
f+5 = 3(s+5)
-------
Substitute for "f" and solve for "s"::
4s+5 = 3s+15
s = 10 (son's age now)
----
f = 4s = 40 (father's age now)
------------------------
Cheers,
Stan H.
---------------